OpenStudy (anonymous):

The salary of a person is increased by 10%, and his expenses also increase by 10%. What is the percentage change in his savings?

OpenStudy (anonymous):

salary be x=x*110/100=11x/10

OpenStudy (anonymous):

expense is x-x/100

OpenStudy (anonymous):

99x/100*110/100

OpenStudy (anonymous):

i think 10% increase

OpenStudy (anonymous):

Apply this(x+y+(xy/100))% where x and y will change sign according to positive or negative variation of given data.

OpenStudy (anonymous):

21

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

Is it decrease of 1%?

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

Is that right or is it 10%?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

@ganeshie8

OpenStudy (cwrw238):

increase of 10%

OpenStudy (anonymous):

how ? explain me

OpenStudy (cwrw238):

if you dont want to you algebra it is valid to take a particular case for example if he earns 100 dollars and has 60 dollars expenses the figures are;- 100 - > 110 60 - . 66 savings 40 --> 44 incraes in savings = 4 / 40 = 10%

OpenStudy (anonymous):

only we can assume expenses right?.

OpenStudy (cwrw238):

yea